OpenStudy (anonymous):

what is the slope of the line that goes through (-10,2) and (-7,8)? I got -6/-17 is this right?

OpenStudy (anonymous):

slop formula is \[\frac{ y2-y1 }{ x2-x1}\]

OpenStudy (anonymous):

I called myself trying to use this and came up with -6/-17

OpenStudy (anonymous):

8-2=6 -7-(-10)=3

OpenStudy (anonymous):

so ur slope would be 6/3 reduce that to 2/1

OpenStudy (anonymous):

I got it! Thank you so much

OpenStudy (anonymous):

No problem :)